6x-3=8x-9, then x=
Answer by TimothyLamb(4379)

---
6x-3 = 8x-9
-2x = -6
x = 3
---
in more detail:
---
6x - 3 = 8x - 9
---
subtract 8x from both sides:
6x - 3 - 8x = 8x - 9 - 8x
-2x - 3 = -9
---
add 3 to both sides:
-2x - 3 + 3 = -9 + 3
-2x = -6
---
divide by -2 both sides:
(-2/-2)x = -6/-2
---
answer:
x = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php